OpenStudy (anonymous):

Select the set of equations that represents the following situation: Mary invested one amount at 8% simple interest, and a second amount at 5% interest, earning $49.16 in one year. If she had switched the amounts, she would have earned $52.37. What were the two amounts? (Points : 3) 8x + 5y = 49.16; 8y + 5x = 52.37 0.08x + 0.05y = 49.16; 0.08y + 0.05x = 52.37 8x + 5y = 52.37; 8x + 5y = 49.16 0.08x + 0.05y = 52.37; 0.08x + 0.05y = 49.16

OpenStudy (anonymous):

8x+5y=52.37;8x+5y=49.16

OpenStudy (anonymous):

noh it sud be 0.08 x + 0.05 y = $ 49.16 0.08 y + 0.05 x = 52.37